• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 07:59:11 on December18,2009

    Entry number 307881

    keyword=Owl shift summary

    Shift started with beam recovery procedure.


    00:13 Run LHRS-26328 RHRS-5832 started with raster 4x4 MCC units and we did spot++.

    00:21 Run 26329 and 5833 started, raster is 6x6 MCC units, we did the spot++.

    00:30 Run 26330 and 5834 started, current is 2 uA, target is LD2, we took Q^2 measurements and posted on halog.

    00:58 Run 26331 and 5835 started, current is 2 uA, target is Aluminum dummy.

    01:13 Run 26332 and 5836 started, current is 2 uA, raster is off, sieve out, optics run.

    01:23 Run 26333 and 5837 started, current is 2 uA, raster is 6x6 MCC units, sieve out, optics run.

    01:37 Run 26334 and 5838 started, 2 uA, 6x6 MCC raster, carbon optics target, left sieve in.

    02:23 Run 26335 and 5839 started, 2 uA, raster off, carbon optics target, left sieve in.

    03:19 Run 26336 and 5840 started, 5 uA, 6x6 MCC raster, LD2 target, left sieve in.

    04:31 Run 26337 and 5841 started, 50 uA, 6x6 MCC raster, Aluminum dummy target, sieve out, VDC off.

    05:06 Run 26338 started, 2 uA, 6x6 MCC raster, we did spot++.

    05:30 Run 26339 and 5842 and parity run 14744 started, 100 uA, 6x6 MCC raster, production.

    06:28 Run 26340 and 5843 and parity run 14745 started, production.

    07:28 Run 26341, 5844 and parity run 14747 started, production.